# Integrating Personality into Digital Humans: A Review of LLM-Driven Approaches for Virtual Reality

### v1

@@ Introduction
Virtual reality (VR) technology has emerged as a powerful medium to create and replicate real-world scenarios in digital spaces, offering unparalleled opportunities for simulation and interaction [23, 53]. Recent advancements in artificial intelligence have further enhanced VR environments by enabling more dynamic scenarios, such as generating text-based descriptions for 360-degree VR scenes [56] and user attention-guidance in virtual tours around real locations using multi-modal models [69]. Complementing these advancements, studies on computer-driven characters with human-like characteristics, referred to as digital humans, highlight their significant role in enhancing user engagement within VR environments by providing realistic facial expressions, responsive gestures, and natural conversational abilities, mimicking humans not only in their visual design but also in their personality, decreasing the gap between virtual and real-world interactions [33, 76].
@@
The potential applications of digital humans in VR environment passes through a variety of fields. In educational settings, the use of virtual pedagogic agents with human-like characteristics demonstrate gains in students’ learning experience and engagement [23], while in therapeutic context, the use of digital humans combined with therapy techniques demonstrates gains in reducing symptoms of borderline personality disorder (BPD) [18]. These possibilities underscore the importance of further research into designing, implementing, and evaluating such agents in VR.
@@
In parallel, the field of natural language processing (NLP) has witnessed transformative advancements with the rise of large language models (LLMs), such as GPT [4] and LLaMA [77] family of models. Trained on vast datasets containing trillions of words, these models demonstrate emergent abilities (i.e., capabilities that arise without explicit training for specific tasks), like the capacity to simulate human personality and emotions in real-time conversations, providing an unprecedented means for human-computer interaction [68, 40, 80]. These capabilities represent a critical step forward in creating digital agents that can think and respond in ways that feel authentically human.
@@
Despite their potential, most applications of these generative models remain confined to chat-based interfaces. Although such systems offer interactive benefits, they do not simulate real-world communication as effectively as VR technology. The integration of NLP-driven digital humans into VR offers a transformative alternative: enabling face-to-face interactions where digital agents can express themselves through both verbal and non-verbal cues, such as facial expressions, gestures, and personality-driven behavior [75, 50]. This multimodal interaction is closer to real human-to-human communication, offering users a more engaging and satisfying experience than chat interfaces.
@@
Furthermore, studies have shown that digital humans personality directly influences how users perceive and interact with the system, impacting key metrics like immersion, trust, and engagement [66, 40]. By embedding these agents with the ability to convey personality traits and human-like expressions, VR can promote deeper connections and more meaningful interactions, paving the way for future advancements in virtual communication.
@@ 1 Digital Humans
Digital humans (DH) have become a focal point of research alongside the rapid expansion of VR technology. These entities are computer-generated representations of human beings, designed to replicate physical appearances, cognitive functions, and emotional responses in digital environments [49, 30]. By simulating human behavior across multiple dimensions, digital humans aim to bridge the gap between artificial systems and natural interactions, unlocking new possibilities for immersive experiences.
Despite their potential, DH often fail to exhibit the nuanced personalities and emotional depth necessary to promote natural and engaging interactions. This absence of relatable human-like traits often results in interactions that feel mechanical and impersonal, making it difficult for users to form meaningful connections with these agents [89, 73]. By designing these agents with more human-like behavior and expressions, these systems can offer greater relatability, which enhances the overall interaction experience and promotes stronger user engagement.
By aligning digital humans behaviors with user expectations and situational needs, these personality-driven agents promote a deeper sense of connection, making interactions more meaningful and enjoyable [66]. Digital humans with empathetic personality demonstrates improvements in learning experiences by creating a supportive environment where users feel understood and motivated, ultimately increasing their willingness to engage and learn [73].
In virtual reality environments, agents with distinct and well-defined personalities, conveyed through both textual and visual features, play a critical role in improving user engagement and immersion. By emulating human-like behaviors and expressions, these agents enable interactions that are both dynamic and natural [73, 58]. Moreover, by embedding distinct personality traits into virtual reality scenarios, they play a pivotal role in enhancing both the realism and the functional effectiveness of digital humans.
@@ 1.1 Personality-driven Agents
Personality conditioning refers to the process of equipping virtual agents with human-like traits and behaviors to enhance the quality of their interactions with users [65]. Early approaches to personality conditioning utilized rule-based systems and decision trees to simulate behavior. These systems mapped specific inputs to predefined outputs, functioning within a limited framework of programmed responses [87]. While innovative for their time, these methods were inherently rigid, lacking the adaptability to handle the unpredictability of human interactions. This rigidity often resulted in repetitive, unnatural conversations that failed to engage users meaningfully or address the complexities of real-world scenarios.
The integration of verbal communication with non-verbal cues has proven essential in enhancing interactions between humans and virtual agents, creating experiences that feel more natural and immersive [6]. By combining elements such as facial expressions, gestures, and tone of voice, these systems enable dynamic and responsive interactions. This multimodal approach not only enriches conversational content but also improves conversational flow, fostering mutual understanding and creating a stronger sense of presence in virtual environments.
@@ 2 Large Language Models (LLMs)
The field of natural language processing (NLP) has undergone rapid transformations since the introduction of the Transformer architecture [81]. This breakthrough has fundamentally reshaped NLP, leading to the proliferation of numerous language models, particularly decoder-based architectures, which have revolutionized text generation and paved the way for large language models (LLMs). These advanced neural networks redefine the boundaries of machine understanding and language generation, leveraging extensive datasets comprising billions or even trillions of words from diverse domains to maximize linguistic coverage and generalization [4].
The unprecedented scale of LLMs, characterized by vast numbers of parameters, enables them to capture intricate linguistic patterns, including semantic relationships, syntactic structures, and long-range dependencies [77, 4]. This capacity allows them to perform complex language-related tasks, such as advanced reasoning and contextual understanding. In particular, LLMs demonstrate emergent abilities, performing tasks that are not explicitly programmed but naturally arise from their scale and the complexity of their training process [85, 37].
The development of large language models generally involves two primary stages: pre-training and fine-tuning. The pre-training phase involves exposing the model to extensive volumes of unlabeled general knowledge text data through a language modeling task, most commonly framed as the next token prediction in a sequence [4]. Then, the resulting model passes through the fine-tuning phase, which refines the model’s capabilities to perform a more specific set of tasks. This process leverages the pre-trained weights, which encode general knowledge about language, morphology, and syntax, and adjusts them using labeled datasets in a supervised learning framework [16]. This step, commonly referred to as Supervised Fine-Tuning (SFT), aligns the model’s performance with the requirements of particular applications, significantly enhancing its ability to tackle domain-specific tasks [52].
While fine-tuning significantly enhances a model’s domain-specific performance, it often requires access to a large amount of labeled data, which presents a substantial barrier, especially when dealing with specialized domains. This challenge has motivated research into alternative approaches to task adaptation, like In-Context Learning (ICL) [47], which allows learning different tasks by textual instructions at inference time using techniques such as zero-shot and few-shot learning [4, 77].
@@ 2.1 Modelling Personality Through LLMs
Building on the concept of emergent abilities in large language models, the task of represent human personality into virtual agents represents an exciting and emergent domain. Personality modeling seeks to create systems that not only respond contextually but also reflect consistent psychological characteristics, enhancing user engagement and creating more lifelike interactions [65].
Research into personality modeling within LLMs has primarily focused on understanding the intrinsic personality traits of these models and developing methods to control or induce specific traits. Studies such as [57] and [67] have analyzed the underlying characteristics of LLMs using established psychological frameworks, such as the MBTI [54] and Big Five [20] personality traits, highlighting the inherent tendencies of LLMs emerged from the training data and model architecture, and exploring how these tendencies might align or deviate from human-like personality expressions.
@@
    By crafting precise instructions that implicitly convey desired personality traits, these methods guide the model to generate responses aligned with specific personality characteristics. For instance, the approach described in [35] uses descriptive prompts that can elicit specific personality traits by framing the context in a way that enables the model to naturally embody the intended personality characteristics.
@@
    This method involves providing the model with a limited number of carefully selected examples that demonstrates the desired personality characteristics, thus improving the ability of the model to generalize these traits in subsequent interactions. An illustrative example is the study [98] where a limited number of carefully crafted examples from the training data (e.g., responses from the IPIP-NEO-120 questionnaire) are provided within the prompt.
@@
    This supervised learning process modifies the pre-trained weights using labeled datasets that reflect the desired personality characteristics, thereby enhancing the model’s ability to consistently exhibit these traits across diverse tasks. An example of this approach is [68] which integrates historical context and behavioral traits into the training process to create historically grounded personas. By refining the model’s internal representations, fine-tuning ensures that the LLM not only understands general language constructs but also reliably manifests the specified personality attributes in its outputs.
@@
Despite significant advancements in personality modeling using large language models (LLMs), much of the research remains centered on text-based chatbot interactions. For example, studies such as [32] and [63] delve into LLM-driven personalities within text or role-playing contexts. Similarly, [39] investigates personality applications in non-player characters (NPCs) within gaming environments. However, these works often neglect the unique challenges of virtual reality, where embodied agents interact with users in immersive, real-time scenarios. Addressing these challenges requires integrating multimodal communication, ensuring responsiveness, and creating an engaging sense of presence, areas that remain underexplored in the current literature.
@@
Virtual reality introduces distinct opportunities by enabling interactions that are richer and more immersive than text-based systems. The use of gestures, facial expressions, and gaze enhances the expression of personality, making interactions more dynamic and engaging. As highlighted in [40], such interactions emulate natural human communication, fostering deeper social connections and user engagement. While these advantages highlight VR’s potential, they also underscore the need for research focused on optimizing personality-driven digital agents for this medium, ensuring that non-verbal communication aligns seamlessly with verbal outputs to create truly lifelike interactions.
@@
The ability to express personality through both text and non-verbal cues, such as gestures and facial expressions, is crucial for enhancing user experience, underscoring the need for further research into seamlessly integrating these expressive modalities into VR environments. While some studies explore multimodal interactions, they often fall short of fully addressing their application in VR. For instance, [5, 12, 83] investigate simulations where agents communicate through speech and gestures or delve into the role of facial expressions, yet they either briefly mention the potential integration of VR or omit it entirely.
@@
Despite the scarcity of research on incorporating LLM-driven personality traits into digital humans in VR, a few studies have begun exploring this convergence. For example, [28] employs multimodal technologies to enhance human-computer interaction, demonstrating the use of LLMs in real-time, open-domain conversations across multiple languages while incorporating speech synthesis and facial expressions. By aligning verbal communication with non-verbal emotional cues, this work illustrates the feasibility of combining LLMs and multimodal systems to create virtual agents with heightened contextual and emotional responsiveness. Such implementations highlight the potential of these technologies to foster more engaging and lifelike interactions in VR environments.
@@ 3 Evaluating Digital Humans Personality Traits
As digital humans become increasingly integrated into daily life, their behavior and impact on user interactions have become areas of growing interest [91, 55]. In the field of psychology, personality traits are defined as enduring patterns of thoughts, feelings, and behaviors that exhibit consistency in time and situations, underscoring individual differences along a set of basic dimensions, reflecting stability and predictability in behavior [15].
@@
The systematic study of personality traits has a long history in psychology, aiming to understand and categorize individual differences [8]. Trait theories provide structured frameworks to explore how people think, feel, and behave in consistent ways over time and across situations. Broad frameworks, like the Big Five model [14], focus on overarching dimensions of personality, organizing personality into five broad dimensions—openness, conscientiousness, extroversion, agreeableness, and neuroticism, capturing fundamental aspects of how individuals engage with the world. More detailed approaches, such as the Sixteen Personality Factors (16PF) [7], break down personality into 16 specific traits, such as emotional stability, social courage, and sensitivity.
@@
Collectively, these methods provide structured tools to explore the complexity of personality, fostering a deeper understanding of the stable characteristics that shape individual differences. Furthermore, research has consistently shown a strong correlation between personality and language use [29, 61, 60, 41]. This relationship has enabled the evaluation of personality traits through text analysis, allowing researchers to assess individual characteristics by examining the frequency and patterns of words and expressions used.
@@
Although foundational theories commonly applied to human beings have inspired efforts to evaluate personality traits in LLMs [35, 27, 88, 82], it remains a complex and nuanced challenge, particularly in digital humans, as it involves a highly subjective context-dependent and non-standardized domain. Existing methods for evaluating LLM personality traits include human evaluations [19, 25], LLM-as-judge approaches [45, 44], and adaptations of self-assessment personality tests originally designed for humans [72, 38].
@@
These approaches have primarily focused on text-based interactions, examining whether LLMs exhibit consistent and stable personality traits [71, 72], exploring methodologies for personality assessment [42, 99], investigating the impact of personality on model safety [92], and tailoring personality traits to specific use cases [46, 31, 82]. These efforts have established a foundation for evaluating personality in text-based agents but fall short when considering the multimodal nature of embodied conversation agents. Due to its complex layers of interaction and connection, we are introduced to challenges that existing evaluation methods, which primarily focus on text-based outputs, are not equipped to address. In that way, significant gaps remain in the standardization and application of multimodal evaluation frameworks.

### v2

# This survey has been expanded and published at Findings of EMNLP 2025. Please cite the version of record: ‘Modeling, Evaluating, and Embodying Personality in LLMs: A Survey” Integrating Personality into Digital Humans: A Review of LLM-Driven Approaches for Virtual Reality

@@ Introduction
Virtual reality (VR) technology has emerged as a powerful medium to create and replicate real-world scenarios in digital spaces, offering unparalleled opportunities for simulation and interaction [pedagogic_agent, mihai_digitaltwin]. Recent advancements in artificial intelligence have further enhanced VR environments by enabling more dynamic scenarios, such as generating text-based descriptions for 360-degree VR scenes [elisasvr] and user attention-guidance in virtual tours around real locations using multi-modal models [paulosvr]. Complementing these advancements, studies on computer-driven characters with human-like characteristics, referred to as digital humans, highlight their significant role in enhancing user engagement within VR environments by providing realistic facial expressions, responsive gestures, and natural conversational abilities, mimicking humans not only in their visual design but also in their personality, decreasing the gap between virtual and real-world interactions [hube2024study, thomas2022investigatinghow].
@@
The potential applications of digital humans in VR environment passes through a variety of fields. In educational settings, the use of virtual pedagogic agents with human-like characteristics demonstrate gains in students’ learning experience and engagement [pedagogic_agent], while in therapeutic context, the use of digital humans combined with therapy techniques demonstrates gains in reducing symptoms of borderline personality disorder (BPD) [Falconer123]. These possibilities underscore the importance of further research into designing, implementing, and evaluating such agents in VR.
@@
In parallel, the field of natural language processing (NLP) has witnessed transformative advancements with the rise of large language models (LLMs), such as GPT [brown2020languagemodelsfewshotlearners] and LLaMA [touvron2023llamaopenefficientfoundation] family of models. Trained on vast datasets containing trillions of words, these models demonstrate emergent abilities (i.e., capabilities that arise without explicit training for specific tasks), like the capacity to simulate human personality and emotions in real-time conversations, providing an unprecedented means for human-computer interaction [shao-etal-2023-character, kroczek2024influencepersonaconversationaltask, tseng-etal-2024-two]. These capabilities represent a critical step forward in creating digital agents that can think and respond in ways that feel authentically human.
@@
Despite their potential, most applications of these generative models remain confined to chat-based interfaces. Although such systems offer interactive benefits, they do not simulate real-world communication as effectively as VR technology. The integration of NLP-driven digital humans into VR offers a transformative alternative: enabling face-to-face interactions where digital agents can express themselves through both verbal and non-verbal cues, such as facial expressions, gestures, and personality-driven behavior [taylor2023rise, maslych2025takeawaysapplyingllmcapabilities]. This multimodal interaction is closer to real human-to-human communication, offering users a more engaging and satisfying experience than chat interfaces.
@@
Furthermore, studies have shown that digital humans personality directly influences how users perceive and interact with the system, impacting key metrics like immersion, trust, and engagement [SACCO2014359, kroczek2024influencepersonaconversationaltask]. By embedding these agents with the ability to convey personality traits and human-like expressions, VR can promote deeper connections and more meaningful interactions, paving the way for future advancements in virtual communication.
@@ 1 Digital Humans
Digital humans (DH) have become a focal point of research alongside the rapid expansion of VR technology. These entities are computer-generated representations of human beings, designed to replicate physical appearances, cognitive functions, and emotional responses in digital environments [lin2024human, HU201164]. By simulating human behavior across multiple dimensions, digital humans aim to bridge the gap between artificial systems and natural interactions, unlocking new possibilities for immersive experiences.
Despite their potential, DH often fail to exhibit the nuanced personalities and emotional depth necessary to promote natural and engaging interactions. This absence of relatable human-like traits often results in interactions that feel mechanical and impersonal, making it difficult for users to form meaningful connections with these agents [WESTER2024100072, sonlu2024effectsembodimentpersonalityexpression]. By designing these agents with more human-like behavior and expressions, these systems can offer greater relatability, which enhances the overall interaction experience and promotes stronger user engagement.
By aligning digital humans behaviors with user expectations and situational needs, these personality-driven agents promote a deeper sense of connection, making interactions more meaningful and enjoyable [SACCO2014359]. Digital humans with empathetic personality demonstrates improvements in learning experiences by creating a supportive environment where users feel understood and motivated, ultimately increasing their willingness to engage and learn [sonlu2024effectsembodimentpersonalityexpression].
In virtual reality environments, agents with distinct and well-defined personalities, conveyed through both textual and visual features, play a critical role in improving user engagement and immersion. By emulating human-like behaviors and expressions, these agents enable interactions that are both dynamic and natural [sonlu2024effectsembodimentpersonalityexpression, pan2024ellmatembodiedllmagentsupporting]. Moreover, by embedding distinct personality traits into virtual reality scenarios, they play a pivotal role in enhancing both the realism and the functional effectiveness of digital humans.
@@ 1.1 Personality-driven Agents
Personality conditioning refers to the process of equipping virtual agents with human-like traits and behaviors to enhance the quality of their interactions with users [SENDRES1995477]. Early approaches to personality conditioning utilized rule-based systems and decision trees to simulate behavior. These systems mapped specific inputs to predefined outputs, functioning within a limited framework of programmed responses [1966]. While innovative for their time, these methods were inherently rigid, lacking the adaptability to handle the unpredictability of human interactions. This rigidity often resulted in repetitive, unnatural conversations that failed to engage users meaningfully or address the complexities of real-world scenarios.
The integration of verbal communication with non-verbal cues has proven essential in enhancing interactions between humans and virtual agents, creating experiences that feel more natural and immersive [Cassell_2001]. By combining elements such as facial expressions, gestures, and tone of voice, these systems enable dynamic and responsive interactions. This multimodal approach not only enriches conversational content but also improves conversational flow, fostering mutual understanding and creating a stronger sense of presence in virtual environments.
@@ 2 Large Language Models (LLMs)
The field of natural language processing (NLP) has undergone rapid transformations since the introduction of the Transformer architecture [vaswani2023attentionneed]. This breakthrough has fundamentally reshaped NLP, leading to the proliferation of numerous language models, particularly decoder-based architectures, which have revolutionized text generation and paved the way for large language models (LLMs). These advanced neural networks redefine the boundaries of machine understanding and language generation, leveraging extensive datasets comprising billions or even trillions of words from diverse domains to maximize linguistic coverage and generalization [brown2020languagemodelsfewshotlearners].
The unprecedented scale of LLMs, characterized by vast numbers of parameters, enables them to capture intricate linguistic patterns, including semantic relationships, syntactic structures, and long-range dependencies [touvron2023llamaopenefficientfoundation, brown2020languagemodelsfewshotlearners]. This capacity allows them to perform complex language-related tasks, such as advanced reasoning and contextual understanding. In particular, LLMs demonstrate emergent abilities, performing tasks that are not explicitly programmed but naturally arise from their scale and the complexity of their training process [wei2022emergentabilitieslargelanguage, kaplan2020scalinglawsneurallanguage].
The development of large language models generally involves two primary stages: pre-training and fine-tuning. The pre-training phase involves exposing the model to extensive volumes of unlabeled general knowledge text data through a language modeling task, most commonly framed as the next token prediction in a sequence [brown2020languagemodelsfewshotlearners]. Then, the resulting model passes through the fine-tuning phase, which refines the model’s capabilities to perform a more specific set of tasks. This process leverages the pre-trained weights, which encode general knowledge about language, morphology, and syntax, and adjusts them using labeled datasets in a supervised learning framework [dong-etal-2024-abilities]. This step, commonly referred to as Supervised Fine-Tuning (SFT), aligns the model’s performance with the requirements of particular applications, significantly enhancing its ability to tackle domain-specific tasks [mecklenburg2024injectingnewknowledgelarge].
While fine-tuning significantly enhances a model’s domain-specific performance, it often requires access to a large amount of labeled data, which presents a substantial barrier, especially when dealing with specialized domains. This challenge has motivated research into alternative approaches to task adaptation, like In-Context Learning (ICL) [li-2023-practical], which allows learning different tasks by textual instructions at inference time using techniques such as zero-shot and few-shot learning [brown2020languagemodelsfewshotlearners, touvron2023llamaopenefficientfoundation].
@@ 2.1 Modelling Personality Through LLMs
Building on the concept of emergent abilities in large language models, the task of represent human personality into virtual agents represents an exciting and emergent domain. Personality modeling seeks to create systems that not only respond contextually but also reflect consistent psychological characteristics, enhancing user engagement and creating more lifelike interactions [SENDRES1995477].
Research into personality modeling within LLMs has primarily focused on understanding the intrinsic personality traits of these models and developing methods to control or induce specific traits. Studies such as [pan2023llmspossesspersonalitymaking] and [serapiogarcía2023personalitytraitslargelanguage] have analyzed the underlying characteristics of LLMs using established psychological frameworks, such as the MBTI [myers1988myers] and Big Five [bigfive] personality traits, highlighting the inherent tendencies of LLMs emerged from the training data and model architecture, and exploring how these tendencies might align or deviate from human-like personality expressions.
@@
    By crafting precise instructions that implicitly convey desired personality traits, these methods guide the model to generate responses aligned with specific personality characteristics. For instance, the approach described in [jiang-etal-2024-personallm] uses descriptive prompts that can elicit specific personality traits by framing the context in a way that enables the model to naturally embody the intended personality characteristics.
@@
    This method involves providing the model with a limited number of carefully selected examples that demonstrates the desired personality characteristics, thus improving the ability of the model to generalize these traits in subsequent interactions. An illustrative example is the study [zhu2024personalityalignmentlargelanguage] where a limited number of carefully crafted examples from the training data (e.g., responses from the IPIP-NEO-120 questionnaire) are provided within the prompt.
@@
    This supervised learning process modifies the pre-trained weights using labeled datasets that reflect the desired personality characteristics, thereby enhancing the model’s ability to consistently exhibit these traits across diverse tasks. An example of this approach is [shao-etal-2023-character] which integrates historical context and behavioral traits into the training process to create historically grounded personas. By refining the model’s internal representations, fine-tuning ensures that the LLM not only understands general language constructs but also reliably manifests the specified personality attributes in its outputs.
@@
Despite significant advancements in personality modeling using large language models (LLMs), much of the research remains centered on text-based chatbot interactions. For example, studies such as [huang2024orcaenhancingroleplayingabilities] and [ran2024capturingmindsjustwords] delve into LLM-driven personalities within text or role-playing contexts. Similarly, [klinkert2024drivinggenerativeagentspersonality] investigates personality applications in non-player characters (NPCs) within gaming environments. However, these works often neglect the unique challenges of virtual reality, where embodied agents interact with users in immersive, real-time scenarios. Addressing these challenges requires integrating multimodal communication, ensuring responsiveness, and creating an engaging sense of presence, areas that remain underexplored in the current literature.
@@
Virtual reality introduces distinct opportunities by enabling interactions that are richer and more immersive than text-based systems. The use of gestures, facial expressions, and gaze enhances the expression of personality, making interactions more dynamic and engaging. As highlighted in [kroczek2024influencepersonaconversationaltask], such interactions emulate natural human communication, fostering deeper social connections and user engagement. While these advantages highlight VR’s potential, they also underscore the need for research focused on optimizing personality-driven digital agents for this medium, ensuring that non-verbal communication aligns seamlessly with verbal outputs to create truly lifelike interactions.
@@
The ability to express personality through both text and non-verbal cues, such as gestures and facial expressions, is crucial for enhancing user experience, underscoring the need for further research into seamlessly integrating these expressive modalities into VR environments. While some studies explore multimodal interactions, they often fall short of fully addressing their application in VR. For instance, [Cai_2024_CVPR, cherakara2023furchatembodiedconversationalagent, 99] investigate simulations where agents communicate through speech and gestures or delve into the role of facial expressions, yet they either briefly mention the potential integration of VR or omit it entirely.
@@
Despite the scarcity of research on incorporating LLM-driven personality traits into digital humans in VR, a few studies have begun exploring this convergence. For example, [hasan_sapienaffective] employs multimodal technologies to enhance human-computer interaction, demonstrating the use of LLMs in real-time, open-domain conversations across multiple languages while incorporating speech synthesis and facial expressions. By aligning verbal communication with non-verbal emotional cues, this work illustrates the feasibility of combining LLMs and multimodal systems to create virtual agents with heightened contextual and emotional responsiveness. Such implementations highlight the potential of these technologies to foster more engaging and lifelike interactions in VR environments.
@@ 3 Evaluating Digital Humans Personality Traits
As digital humans become increasingly integrated into daily life, their behavior and impact on user interactions have become areas of growing interest [yang2024makesmodellowempathywarmth, noever2023aitexttobehaviorstudysteerability]. In the field of psychology, personality traits are defined as enduring patterns of thoughts, feelings, and behaviors that exhibit consistency in time and situations, underscoring individual differences along a set of basic dimensions, reflecting stability and predictability in behavior [diener2019personality].
@@
The systematic study of personality traits has a long history in psychology, aiming to understand and categorize individual differences [cattell1977scientific]. Trait theories provide structured frameworks to explore how people think, feel, and behave in consistent ways over time and across situations. Broad frameworks, like the Big Five model [de2000big], focus on overarching dimensions of personality, organizing personality into five broad dimensions—openness, conscientiousness, extroversion, agreeableness, and neuroticism, capturing fundamental aspects of how individuals engage with the world. More detailed approaches, such as the Sixteen Personality Factors (16PF) [cattell2008sixteen], break down personality into 16 specific traits, such as emotional stability, social courage, and sensitivity.
@@
Collectively, these methods provide structured tools to explore the complexity of personality, fostering a deeper understanding of the stable characteristics that shape individual differences. Furthermore, research has consistently shown a strong correlation between personality and language use [hirsh2009personality, pennebaker1999linguistic, pennebaker2001patterns, lee2007relations]. This relationship has enabled the evaluation of personality traits through text analysis, allowing researchers to assess individual characteristics by examining the frequency and patterns of words and expressions used.
@@
Although foundational theories commonly applied to human beings have inspired efforts to evaluate personality traits in LLMs [jiang-etal-2024-personallm, hagendorff2023machine, wen2024selfassessmentexhibitionrecognitionreview, vu2025psychadapteradaptingllmtransformers], it remains a complex and nuanced challenge, particularly in digital humans, as it involves a highly subjective context-dependent and non-standardized domain. Existing methods for evaluating LLM personality traits include human evaluations [feng2024sampleefficienthumanevaluationlarge, guzman-etal-2015-humans], LLM-as-judge approaches [li2024llmsasjudgescomprehensivesurveyllmbased, li2025generationjudgmentopportunitieschallenges], and adaptations of self-assessment personality tests originally designed for humans [song2023largelanguagemodelsdeveloped, karra2023estimatingpersonalitywhiteboxlanguage].
@@
These approaches have primarily focused on text-based interactions, examining whether LLMs exhibit consistent and stable personality traits [song2024identifyingmultiplepersonalitieslarge, song2023largelanguagemodelsdeveloped], exploring methodologies for personality assessment [lee2024llmsdistinctconsistentpersonality, zou2024llmselfreportevaluatingvalidity], investigating the impact of personality on model safety [zhang2024betterangelsmachinepersonality], and tailoring personality traits to specific use cases [li2024big5chatshapingllmpersonalities, huang2024designingllmagentspersonalitiespsychometric, vu2025psychadapteradaptingllmtransformers]. These efforts have established a foundation for evaluating personality in text-based agents but fall short when considering the multimodal nature of embodied conversation agents. Due to its complex layers of interaction and connection, we are introduced to challenges that existing evaluation methods, which primarily focus on text-based outputs, are not equipped to address. In that way, significant gaps remain in the standardization and application of multimodal evaluation frameworks.